OpenStudy (anonymous):

EVERYBODY LOOK HERE!!!!! what is -6.4 as a fraction

OpenStudy (anonymous):

Gracias Senor I am forever grateful!!!! U get fre medal and fan and...thats all i can give:(

OpenStudy (karatechopper):

-6 2/5

OpenStudy (karatechopper):

and i am a Senorita

OpenStudy (anonymous):

gracias senorita ;)

OpenStudy (anonymous):

-6.4 = -(6+0.4) = -(6+2/5) = -32/5

OpenStudy (turingtest):

ok, you have my attention... rewrite this as\[-\frac{6.4}{1}\]now we want to get rid of the 0.4 on top, so we need to multiply it by something to make it go away. Whatever we multiply the top by though, we have to multiply the bottom by as well. A little thought and you will notice that \[6.4\times5=32\]so we will multiply top and bottom of our fraction by 5\[-\frac{6.4}{1}\times\frac{5}{5}=-\frac{32}{5}\]try to understand the process I used so you don't have to keep asking the same type of question.

OpenStudy (karatechopper):

:D

OpenStudy (anonymous):

ALL OF YOU HAVE MY ETERNAL LOVE!!!

OpenStudy (akshay_budhkar):

nothing wrong is what turing says.. its more than perfect.. just to add my toppings on that beautiful cake, i will multiply it with 10 instead of 5 then divide it with 2 :D Find that simpler and faster :D

OpenStudy (turingtest):

true^ smarter, and therefor easier way :)

OpenStudy (akshay_budhkar):

:D

OpenStudy (anonymous):

Turing's was good because he saw that 5 was needed to multiply 6.4 in order to produce an integer. In fact, and integer multiple of 5 will do, picking 5 shows mad skillz

OpenStudy (anonymous):

so how woud u do 0.48

OpenStudy (anonymous):

any*

OpenStudy (turingtest):

Let's do this one akshay's way:\[0.48=\frac{0.48}{1}\]the top will obviously be an integer if we multiply it by 100, because\[0.48\times100=48\]but that means we need to do the same to the bottom as well. So what do we get by multiplying the top and bottom by 100 ?

OpenStudy (anonymous):

um...0.480/100?

OpenStudy (turingtest):

^the numerator is wrong.\[0.48\times100=48\]so\[0.48=\frac{0.48}{1}\times\frac{100}{100}=?\]

OpenStudy (anonymous):

karatechopper help me plzz!!

OpenStudy (anonymous):

ok so 48/100

OpenStudy (turingtest):

remember that whenever we multiply a decimal by some power of ten, we move the decimal point to the right by the number of zeros after the one 100 has two zero's, so\[0.48\times100\]will work like so now we have\[0.48=\frac{48}{100}\]notice that both the top and bottom are even, so we can divide them both by 2 (this was akshay's method) so what do we get by dividing the top and bottom of our fraction by 2?